Find the sum of the fcllowing geometric series.
8 + 12 +..... + 60.75

Answer by venugopalramana(3286)

Find the sum of the fcllowing geometric series.
8 + 12 +..... + 60.75
A=8
R=12/8=1.5
TN=AR^(N-1)=8*(3/2)^(N-1)=60.75
N=6 BY INSPECTION SINCE 8*3*3*3*3*3/(2*2*2*2*2)=60.75
SN=S6=A*{R^N - 1)/(R-1)=(60.75*1.5-8)/0.5=166.25